# Covariant Quantum Error-Correcting Codes with Metrological Entanglement Advantage

### v1

@@ Abstract
If $\theta$ is generated from $e^{-i\theta\hat{Q}^{z}}$, then without entanglement, the QFI is upper bounded by the standard quantum limit, scaling as $\mathfrak{F}\sim N$. Achieving $\mathfrak{F}\sim N^{\gamma}$ for $1<\gamma\leq 2$ indicates a metrological entanglement advantage, with $\gamma=2$ as the Heisenberg limit. Since $\mathfrak{F}(\psi_{\theta};\partial_{\theta}\psi_{\theta})=4(\langle\psi_{\theta}|(\hat{Q}^{z})^{2}|\psi_{\theta}\rangle-\langle\psi_{\theta}|\hat{Q}^{z}|\psi_{\theta}\rangle^{2})$ for a pure state, our code can have an entanglement advantage if it contains a logical state whose variance of $\hat{Q}^{z}$ scales super-linearly with $N$. This is possible with the probe state $|\psi_{0}\rangle\propto|J,M\rangle+|J,-M\rangle$, whose QFI is $\mathfrak{F}(\psi_{\theta})=4M^{2}\!\sim\!N^{2b}$ for $M\!\sim\!N^{b}$ and $b>\frac{1}{2}$.
While a GHZ(-like) state $|\psi^{\prime}\rangle=(|J,J\rangle+|J,-J\rangle)/\sqrt{2}$ can achieve the Heisenberg limit [42, 43, 44], its QFI becomes zero even if only one qudit is erased (see App. S7). Various other probe states also provide an entanglement advantage, e.g., spin-squeezed [45, 46, 47, 48, 49], Dicke [44, 50], and scrambled [51] states. Additionally, various studies have employed different methods to assess the the effect noise in metrology [52, 53, 54, 55, 56, 57, 58]. Differently, we use AQECC inaccuracy to rigorously bound the QFI loss under certain noise, showing that entanglement advantage persists. Notably, the QFI protection here is achieved without repeated quantum error correction [10], differing from the typical error-corrected metrology scenarios [1, 2, 3, 4, 5, 6, 7, 8, 11]. In a recent work [59], a non-maximally symmetric state also achieves metrological entanglement advantage without error correction.
@@
$\mathfrak{F}(\psi_{\theta};\partial_{\theta}\psi_{\theta})-\mathfrak{F}(\mathcal{M}(\psi_{\theta});\mathcal{M}(\partial_{\theta}\psi))\leq 16\epsilon\|\partial_{\theta}\psi_{\theta}\|_{1}\|\partial_{\theta}\psi_{\theta}\|_{\text{op}}~{}.$ (6)
@@
Since $\|\mathcal{E}-\mathcal{F}\|_{\diamond}\leq 2D_{P}(\mathcal{E},\mathcal{F})$ (see App. S1), taking $\mathcal{M}=\mathcal{R}\circ\mathcal{N}$, we have $\|\mathcal{M}-\mathrm{id}\|_{\diamond}\leq 2\epsilon$ bounded by the AQECC inaccuracy. In fact, from the data processing inequality [60, 10], we have $\mathfrak{F}(\mathcal{R}\circ\mathcal{N}(\rho_{\theta});\mathcal{R}\circ\mathcal{N}(\partial_{\theta}\rho_{\theta}))\leq\mathfrak{F}(\mathcal{N}(\rho_{\theta});\mathcal{N}(\partial_{\theta}\rho_{\theta}))$. That is, the protection of QFI can already be achieved by encoding alone.
@@
For metrology purposes, $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle,|J\!,\!-M\rangle\}$ suffices to host the probe state $|\psi_{0}\rangle=(|J\!,\!M\rangle+|J\!,\!-M\rangle)/\sqrt{2}$, which has $\|\partial_{\theta}\psi_{\theta}\|_{1}=\|\partial_{\theta}\psi_{\theta}\|_{\text{op}}=M$. Therefore, we have $\mathfrak{F}(\mathcal{N}(\psi_{\theta});\mathcal{N}(\partial_{\theta}\psi_{\theta}))\geq(1-4\epsilon)4M^{2}$, and a protected entanglement advantage is possible if $\epsilon\rightarrow 0$ when $N\rightarrow\infty$ while having $M\sim N^{b}$ with $b>1/2$.
@@
Now considering the $d$-local noise, for $J\sim N^{a}$, $M\sim N^{b}$, and $d\sim c\log_{2}(N)$, Theorem 1 gives $\epsilon=\tilde{O}(N^{(2c\log_{2}(2s+1)+b-a)/2})$. So for $a>b>1/2$, the noisy probe state can maintain an entanglement advantage $\mathfrak{F}(\mathcal{N}(\psi_{\theta}))=\Theta(N^{2b})$ if $c<(a-b)/(4\log_{2}(2s+1))$ (illustrated in Fig. 1(a)). For heralded $d$-local erasure errors with $d\sim N^{c}$, Theorem 2 gives $\epsilon=O(N^{b+c/2-1})$. The partially-erased probe state still has an entanglement advantage $\mathfrak{F}(\mathcal{N}(\psi_{\theta}))=\Theta(N^{2b})$ with $21>b>1/$ 2 and $\epsilon\rightarrow 0$ provided $c<2(1-b)$ and $c\leq b$ (illustrated in Fig. 1(b)).
@@
Figure 2: QFI loss $\Delta\mathfrak{F}$ of the probe state $|\psi\rangle\propto|J,M\rangle+|J,-M\rangle$ ($J=sN$) for (a) $M\sim J^{\frac{3}{5}}$, (b) $M\sim J^{\frac{2}{3}}$, and (c) $M\sim J^{\frac{3}{4}}$, after erasing $d=j_{1}/s$ qudits. We calculate $\Delta\mathfrak{F}$ by picking values of $J$ and $j_{1}$, giving us results for any spin-$s$ and corresponding $N\!=\!J/s$ and $d=j_{1}/s$. Numerical results show $\Delta\mathfrak{F}/(4M^{2})\sim J^{2b+c-2}$ for the parameters in the red-shaded region of Fig. 1(b).
@@
To assess our QFI-loss bound, we numerically calculate $\Delta\mathfrak{F}:=\mathfrak{F}(\psi_{\theta})-\mathfrak{F}(\mathcal{N}(\psi_{\theta}))$ for $J=sN$, where the channel is $\mathcal{N}(\psi_{\theta})=\text{Tr}_{\alpha}[\psi_{\theta}]$ and $\alpha$ contains $d$ spins (so $j_{1}=sd$). In Fig. 2, we plot $\Delta\mathfrak{F}/(4M^{2})$ for various $b$ and $c$ in the red-shaded region of Fig. 1(b). The numerical calculations are detailed in App. S7. Applying the QFI-loss bound Eq. (6), we have $\Delta\mathfrak{F}/(4M^{2})\leq 4\epsilon$. We observe $\Delta\mathfrak{F}/M^{2}\sim J^{2b+c-2}$, which is indeed asymptotically upper-bounded by $J^{b+c/2-1}\sim N^{b+c/2-1}$ Though the bound is not tight, it could be due to the Fuchs-van de Graaf inequality [61] not being saturated asymptotically. However, within the red region in Fig. 1(b), $\mathfrak{F}(\psi_{\theta})$ and $\mathfrak{F}(\mathcal{N}(\psi_{\theta}))$ both scale as $N^{2b}$, both indicating a metrological entanglement advantage.
@@
In App. S7, we present a local measurement that saturates the quantum Cramér-Rao bound for the ideal probe state $|\psi_{\theta}\rangle$. For the partially-erased probe state, a global measurement can achieve the QFI scaling, while a local measurement can surpass the standard limit in certain parameter regimes, despite not saturating the quantum Cramér-Rao bound.
@@
Discussion.—We establish code-inaccuracy bounds for a family of covariant approximate quantum error-correcting codes whose codewords house irreducible representations of $SU(2)$. We show that a particular (entangled) codeword can be used for quantum sensing in a way that outperforms classical strategies, even in the presence of certain noise.
Ref. [62] shows that an $N$-qubit Dicke state $|J=\frac{N}{2}\!,\!M\rangle$ can be prepared using a quantum circuit with a number of gates polynomial in $N$ and $M$. One can then prepare the probe state $|\psi_{0}\rangle$ using an ancilla qubit and a controlled version of this circuit. Exploring whether this method can be extended to qudits would be valuable.
@@
We have proposed some measurement schemes for the ideal and partially-erased probe states in App. S7. It is worth investigating whether simpler measurements can saturate the Cramér-Rao bound for our noisy state [63]. While the recovery channel [18, 64] cannot increase quantum Fisher information [60, 10], it may aid in devising a measurement protocol for any noisy probe.
@@
With modifications, our approximate quantum error-correcting code results can be applied to other $SU(2)$ code words, like the quantum many-body scar states in the spin-1 XY-Dzyaloshinskii-Moriya model [65, 66, 67] (see App. S8 for details). Accordingly, the $SU(2)$-scar probe states could be noise-resilient, complementing the results using scar states for sensing [67, 68, 69]. For states with approximate $SU(2)$ symmetry [24, 25], like the scars in a deformed PXP model [70, 71, 72, 73, 74], some exact codes [24, 25], or approximately prepared states [75], further study is warranted to determine if they also provide a noise-resilient entanglement advantage.
@@
Our result suggests a new way of finding approximate quantum error-correcting codes with continuous transversal logical gates by using symmetry irreps, complementing existing methods [32, 14, 33, 29, 31]. Here, one selects a subset of the $SU(2)$-irrep basis to form an approximate quantum error-correcting code, reducing the symmetry to $U(1)$ and giving rise to continuous transversal logical gates. This raises the question of whether one can also devise approximate quantum error-correcting codes by using irreps from other continuous symmetries (say, $SU(3)$), which could potentially yield transversal gates generated by multiple (possibly non-commuting) generators and an entanglement advantage for multi-parameter sensing.
@@ Acknowledgements.
We thank Xiaozhen Fu, Shubham Jain, Zhi Li, Zhenning Liu, Sean R. Muleady, and Sisi Zhou for valuable discussions and feedback. C.-J. Lin acknowledges the support from the National Science Foundation (QLCI grant OMA-2120757). Z.-W.L. is supported in part by a startup funding from YMSC, Tsinghua University, and NSFC under Grant No. 12475023. A.V.G. was supported in part by AFOSR MURI, DARPA SAVaNT ADVENT, NSF QLCI (award No. OMA-2120757), DoE ASCR Accelerated Research in Quantum Computing program (award No. DE-SC0020312), the DoE ASCR Quantum Testbed Pathfinder program (awards No. DE-SC0019040 and No. DE-SC0024220), and the NSF STAQ program. Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator.
@@
This Supplemental Material is organized as follows. In Sec. S1, we list the definitions of various quantities used throughout the paper. We then present the proofs of Lemma 1 (Sec. S2), Lemma 2 (Sec. S3), Theorem 1 & Corollary 1 (Sec. S4), Lemma 3 (Sec. S5), and Theorem 2 (Sec. S6). In particular, Lemma 1 pertains to the “off-diagonal” part of the Knill-Laflamme conditions, while Lemma 2 bounds the “diagonal” part. Theorem 1 and Corollary 1 give a bound on the code inaccuracy against a generic $d$-local noise channel. Lemma 3 pertains to the fidelity between the reduced density matrices from code words of different magnetic quantum numbers. Theorem 2 describes an upper bound on the code inaccuracy against a heralded $d$-local erasure error. In Sec. S7, we provide the details behind the numerical calculation of the quantum Fisher information and behind the measurement protocols described in the main text. Finally, in Sec. S8, we show that the quantum many-body scar states in the spin-1 XY-Dzyaloshinskii-Moriya model can also form an covariant AQECC.
@@
(Trace norm) $\|\rho\|_{1}:=\text{Tr}\sqrt{\rho^{\dagger}\rho}$ ,
which induces the trace distance between two density matrices $\rho$ and $\sigma$ as $\|\rho-\sigma\|_{1}$. The trace distance then induces a distance measure between quantum channels $\mathcal{E}$ and $\mathcal{F}$:
@@
(Fidelity) $f(\rho,\sigma):=\|\sqrt{\sigma}\sqrt{\rho}\|_{1}=\text{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}$ ,
@@
(Purified channel distance) $D_{P}(\mathcal{E},\mathcal{F}):=\sqrt{1-F(\mathcal{E},\mathcal{F})^{2}}$
@@
(Bures channel distance) $D_{B}(\mathcal{E},\mathcal{F}):=\sqrt{1-F(\mathcal{E},\mathcal{F})}$
@@
Lemma 2: Assume $\hat{F}$ is a $d$-local operator, then $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|\leq dq_{0}\|\hat{F}\|_{op}C(n,m)$, where $C(n,m):=\sum_{M=n}^{m-1}[c^{+}_{M}]^{-1}$ and $q_{0}$ is a constant. In particular, for $0\leq b<a\leq 1$, if we take $J\sim N^{a}$ and $|m-n|\sim N^{b}$, then we have $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|=\|\hat{F}\|_{op}O(dN^{b-a})$.
@@
We can obtain a bound on $C(n,m)$ as follows. Consider the function $g(x):=[(J-x)(J+x+1)]^{-1/2}$ in the domain $x\in(-J-1,J)$. It is easy to see that the minimum of $g(x)$ is at $x^{*}=-1/2$, and $g(x)$ is a decreasing function from $x=-J-1$ to $x=x^{*}$ and an increasing function from $x=x^{*}$ to $x=J$. It is also easy to check that $g(x)=g(2x^{*}-x)$. We first consider the case $n<0<m$. We separate $C(n,m)$ into two pieces, $C(n,m)=C_{1}(n)+C_{2}(m)$, where $C_{1}(n):=\sum_{\ell=n}^{\ell=-1}[c^{+}_{\ell}]^{-1}$ and $C_{2}(m):=\sum_{\ell=0}^{m-1}[c^{+}_{\ell}]^{-1}$. These two pieces can be bounded as
[…]
@@
[…]
@@
$|\langle J,{m}|\hat{F}|J,{m}\rangle-\langle J,n|\hat{F}|J,n\rangle|=\|\hat{F}\|_{op}O(dN^{b-a})~{}.$ (S12)
@@
Theorem 1: Consider the code space $\mathfrak{C}=span\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J\sim N^{a}$, $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\sim N^{b}$, and $b<a\leq 1$. $\mathfrak{C}$ forms an $(\!(N,O(b\log_{2}N))\!)$ AQECC against a $d$-local noise channel with $\epsilon(d)=O(\sqrt{d}(2s+1)^{d}N^{(b-a)/2})$ if we take $\Delta\geq 2sd+1$.

### v2

@@ Abstract
If $\theta$ is generated from $e^{-i\theta\hat{Q}^{z}}$, then without entanglement, the QFI is upper bounded by the standard quantum limit, scaling as $\mathfrak{F}\!\sim\!N$. Achieving $\mathfrak{F}\!\sim\!N^{\gamma}$ for $1\!<\!\gamma\!\leq\!2$ indicates a metrological entanglement advantage, with $\gamma\!=\!2$ as the Heisenberg limit. Since $\mathfrak{F}(\psi_{\theta};\partial_{\theta}\psi_{\theta})=4[\langle\psi_{\theta}|(\hat{Q}^{z})^{2}|\psi_{\theta}\rangle-\langle\psi_{\theta}|\hat{Q}^{z}|\psi_{\theta}\rangle^{2}]$ for a pure state, our code can have an entanglement advantage if it contains a logical state whose variance of $\hat{Q}^{z}$ scales superlinearly with $N$. This is possible with the probe state $|\psi_{0}\rangle\propto|J,M\rangle+|J,-M\rangle$, whose QFI is $\mathfrak{F}(\psi_{\theta})=4M^{2}\!\sim\!N^{2b}$ for $M\!\sim\!N^{b}$ and $b>\frac{1}{2}$.
While a GHZ(-like) state $|\psi^{\prime}\rangle=(|J,J\rangle+|J,-J\rangle)/\sqrt{2}$ can achieve the Heisenberg limit [43, 44, 45], its QFI becomes zero even if only one qudit is erased (see Supplemental Material [40] S8). Various other probe states also provide an entanglement advantage, e.g., spin-squeezed [46, 47, 48, 49, 50], Dicke [45, 51, 52], and scrambled [53] states. Additionally, various studies have employed different methods to assess the effect noise in metrology [54, 55, 56, 57, 58, 59, 60]. Differently, we use AQECC inaccuracy to rigorously bound the QFI loss under certain noise, showing that entanglement advantage persists. Notably, the QFI protection here is achieved without repeated quantum error correction [10], differing from the typical error-corrected metrology scenarios [1, 2, 3, 4, 5, 6, 7, 8, 11]. In a recent work [61], a nonmaximally symmetric state also achieves metrological entanglement advantage without error correction.
@@
$\mathfrak{F}(\psi_{\theta};\partial_{\theta}\psi_{\theta})-\mathfrak{F}[\mathcal{M}(\psi_{\theta});\mathcal{M}(\partial_{\theta}\psi)]\leq 16\epsilon\|\partial_{\theta}\psi_{\theta}\|_{1}\|\partial_{\theta}\psi_{\theta}\|_{\text{op}}~.$ (7)
@@
Since $\|\mathcal{E}-\mathcal{F}\|_{\diamond}\leq 2D_{P}(\mathcal{E},\mathcal{F})$ (see Supplemental Material S1 [40]), taking $\mathcal{M}=\mathcal{R}\circ\mathcal{N}$, we have $\|\mathcal{M}-\mathrm{id}\|_{\diamond}\leq 2\epsilon$ bounded by the AQECC inaccuracy. In fact, from the data processing inequality [62, 10], we have $\mathfrak{F}[\mathcal{R}\circ\mathcal{N}(\rho_{\theta});\mathcal{R}\circ\mathcal{N}(\partial_{\theta}\rho_{\theta})]\leq\mathfrak{F}[\mathcal{N}(\rho_{\theta});\mathcal{N}(\partial_{\theta}\rho_{\theta})]$. That is, the protection of QFI can already be achieved by encoding alone.
@@
For metrology purposes, $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle,|J\!,\!-M\rangle\}$ suffices to host the probe state $|\psi_{0}\rangle=(|J\!,\!M\rangle+|J\!,\!-M\rangle)/\sqrt{2}$, which has $\|\partial_{\theta}\psi_{\theta}\|_{1}=\|\partial_{\theta}\psi_{\theta}\|_{\text{op}}=M$. Therefore, we have $\mathfrak{F}[\mathcal{N}(\psi_{\theta});\mathcal{N}(\partial_{\theta}\psi_{\theta})]\geq(1-4\epsilon)4M^{2}$, and a protected entanglement advantage is possible if $\epsilon\rightarrow 0$ when $N\rightarrow\infty$ while having $M\sim N^{b}$ with $b>1/2$.
@@
Now considering the $d$-local noise, for $J\!\sim\!N^{a}$, $M\!\sim\!N^{b}$, and $d\!\sim\!c\log_{2}(N)$, Theorem 1 gives $\epsilon=\tilde{O}[N^{-(a-b-2c\log_{2}(2s+1))/2}]$. So for $a>b>1/2$, the noisy probe state can maintain an entanglement advantage $\mathfrak{F}[\mathcal{N}(\psi_{\theta})]=\Theta(N^{2b})$ if $c<(a-b)/[4\log_{2}(2s+1)]$ [illustrated in Fig. 1(a)]. For heralded $d$-local erasure errors with $d\sim N^{c}$, Theorem 2 gives $\epsilon=O(N^{b+c/2-1})$. The partially-erased probe state still has an entanglement advantage $\mathfrak{F}[\mathcal{N}(\psi_{\theta})]=\Theta(N^{2b})$ with $1>b>1/2$ and $\epsilon\rightarrow 0$ provided c<2(1b)c<2(1-b) and $c\leq b$ [illustrated in Fig. 1(b)].
@@
Figure 2: QFI loss $\Delta\mathfrak{F}$ of the probe state $|\psi\rangle\propto|J,M\rangle+|J,-M\rangle$ ($J=sN$) for (a) $M\sim J^{\frac{3}{5}}$, (b) $M\sim J^{\frac{2}{3}}$, and (c) $M\sim J^{\frac{3}{4}}$, after erasing $d=j_{1}/s$ qudits. We calculate $\Delta\mathfrak{F}$ by picking values of $J$ and $j_{1}$, giving us results for any spin $s$ and corresponding N=J/sN\!=\!J/s and $d=j_{1}/s$. Numerical results show $\Delta\mathfrak{F}/(4M^{2})\sim J^{2b+c-2}$ for the parameters in the red-shaded region of Fig. 1(b).
@@
To assess our QFI-loss bound, we numerically calculate $\Delta\mathfrak{F}:=\mathfrak{F}(\psi_{\theta})-\mathfrak{F}[\mathcal{N}(\psi_{\theta})]$ for $J=sN$, where the channel is $\mathcal{N}(\psi_{\theta})=\mathrm{Tr}_{\alpha}[\psi_{\theta}]$ and $\alpha$ contains $d$ spins (so $j_{1}=sd$). In Fig. 2, we plot $\Delta\mathfrak{F}/(4M^{2})$ for various $b$ and $c$ in the red-shaded region of Fig. 1(b). The numerical calculations are detailed in Supplemental Material S8 [40]. Applying the QFI-loss bound Eq. (7), we have $\Delta\mathfrak{F}/(4M^{2})\leq 4\epsilon$. We observe that $\Delta\mathfrak{F}/M^{2}\sim J^{2b+c-2}$, which is indeed asymptotically upper bounded by $J^{b+c/2-1}\sim N^{b+c/2-1}$ . Though the bound is not tight, it could be due to the Fuchs-van de Graaf inequality [63] not being saturated asymptotically. However, within the red region in Fig. 1(b), $\mathfrak{F}(\psi_{\theta})$ and $\mathfrak{F}[\mathcal{N}(\psi_{\theta})]$ both scale as $N^{2b}$, both indicating a metrological entanglement advantage.
@@
In Supplemental Material S8 [40], we present a local measurement that saturates the quantum Cramér-Rao bound for the ideal probe state $|\psi_{\theta}\rangle$. For the partially erased probe state, a global measurement can achieve the QFI scaling, while a local measurement can surpass the standard limit in certain parameter regimes, despite not saturating the quantum Cramér-Rao bound.
@@
Discussion—We establish code-inaccuracy bounds for a family of covariant approximate quantum error-correcting codes whose codewords house irreducible representations of SU(2). We show that a particular (entangled) codeword can be used for quantum sensing in a way that outperforms classical strategies, even in the presence of certain noise.
References [64, 65, 66, 67, 68] show protocols to prepare $|J\!,\!M\rangle$ states for $J=sN$. One can then prepare the probe state $|\psi_{0}\rangle$ using an ancilla qubit and a controlled version of the circuit that prepares $|J\!,\!M\rangle$.
@@
We have proposed some measurement schemes for the ideal and partially-erased probe states in Supplemental Material S8 [40]. It is worth investigating whether simpler measurements can saturate the Cramér-Rao bound for our noisy state [69]. While the recovery channel [18, 70] cannot increase quantum Fisher information [62, 10], it may aid in devising a measurement protocol for any noisy probe. Refs. [18, 71] proposed a near-optimal recovery channel for AQECCs, which we outline in Supplemental Material S9 [40]. Although an efficient circuit implementation is not guaranteed, it would be interesting to explore whether the structure of the $J=sN$ code, such as its total permutation symmetry, can be leveraged for this purpose.
@@
With modifications, our approximate quantum error-correcting code results can be applied to other SU(2) code words, like the quantum many-body scar states in the spin-1 XY-Dzyaloshinskii-Moriya model [72, 73, 74] (see Supplemental Material S10 [40] for details). Accordingly, the SU(2)-scar probe states could be noise-resilient, complementing the results using scar states for sensing [74, 75, 76]. For states with approximate SU(2) symmetry [24, 25], like the scars in a deformed PXP model [77, 78, 79, 80, 81], some exact codes [24, 25], or approximately prepared states [82], further study is warranted to determine if they also provide a noise resilient entanglement advantage.
@@
Our result suggests a new way of finding approximate quantum error-correcting codes with continuous transversal logical gates by using symmetry irreps, complementing existing methods [32, 14, 33, 29, 31]. Here, one selects a subset of the SU(2)-irrep basis to form an approximate quantum error-correcting code, reducing the symmetry to U(1) and giving rise to continuous transversal logical gates. This raises the question of whether one can also devise approximate quantum error-correcting codes by using irreps from other continuous symmetries [say, SU(3)], which could potentially yield transversal gates generated by multiple (possibly noncommuting) generators and an entanglement advantage for multiparameter sensing.
@@ Acknowledgements.
We thank Xiaozhen Fu, Shubham Jain, Zhi Li, Zhenning Liu, Sean R. Muleady, Nicole Yunger Halpern, and Sisi Zhou for valuable discussions and feedback. C.-J. Lin acknowledges the support from the National Science Foundation (QLCI Grant No. OMA-2120757). Z.-W.L. is supported in part by a startup funding from YMSC, Tsinghua University, Dushi Program, and NSFC under Grant No. 12475023. A.V.G. was supported in part by AFOSR MURI, DARPA SAVaNT ADVENT, NSF QLCI (Award No. OMA-2120757), the DoE ASCR Quantum Testbed Pathfinder program (Awards No. DE-SC0019040 and No. DE-SC0024220), NSF STAQ program, ARL (W911NF-24-2-0107), and NQVL:QSTD:Pilot:FTL. A.V.G. also acknowledges support from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (QSA) and from the U.S. Department of Energy, Office of Science, Accelerated Research in Quantum Computing, Fundamental Algorithmic Research toward Quantum Utility (FAR-Qu).
[…]
@@
This Supplemental Material is organized as follows. In Sec. S1, we list the definitions of various quantities used throughout the paper. We then present the proofs of Lemma 1 (Sec. S2), Lemma 2 (Sec. S3), Theorem 1 & Corollary 1 (Sec. S4), Corollary 2 (Sec. S5), Lemma 3 (Sec. S6), and Theorem 2 (Sec. S7). In particular, Lemma 1 pertains to the “off-diagonal” part of the Knill-Laflamme conditions, while Lemma 2 bounds the “diagonal” part. Theorem 1 and Corollary 1 give a bound on the code inaccuracy against a generic $d$-local noise channel. Corollary 2 gives conditions and an explicit bound on the code inaccuracy under independent and identically distributed (i.i.d.) noise. Lemma 3 pertains to the fidelity between the reduced density matrices from code words of different magnetic quantum numbers. Theorem 2 describes an upper bound on the code inaccuracy against a heralded $d$-local erasure error. In Sec. S8, we provide the details behind the numerical calculation of the quantum Fisher information and behind the measurement protocols described in the main text. We discuss a construction of a recovery channel in Sec. S9. Finally, in Sec. S10, we show that the quantum many-body scar states in the spin-1 XY-Dzyaloshinskii-Moriya model can also form an covariant AQECC.
@@
(Operator norm) $\|\hat{F}\|_{\text{op}}:=\sup_{|\psi\rangle}\{\|\hat{F}|\psi\rangle\|:\||\psi\rangle\|\!=\!1\}$
is the operator norm, where $\||\psi\rangle\|:=\sqrt{\langle\psi|\psi\rangle}$ is the vector norm.
Definition 2.
(Trace norm) $\|\rho\|_{1}:=\mathrm{Tr}\sqrt{\rho^{\dagger}\rho}$ ,
which induces the trace distance between two density matrices $\rho$ and $\Gamma$ as $\|\rho-\Gamma\|_{1}$. The trace distance then induces a distance measure between quantum channels $\mathcal{E}$ and $\mathcal{F}$:
@@
(Fidelity) $f(\rho,\sigma):=\|\sqrt{\sigma}\sqrt{\rho}\|_{1}=\mathrm{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}$ ,
@@
(Purified channel distance) $D_{P}(\mathcal{E},\mathcal{F}):=\sqrt{1-F(\mathcal{E},\mathcal{F})^{2}}$ .
@@
(Bures channel distance) $D_{B}(\mathcal{E},\mathcal{F}):=\sqrt{1-F(\mathcal{E},\mathcal{F})}$ .
@@
Lemma 2: Assume $\hat{F}$ is a $d$-local operator, then $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|\leq dq_{0}\|\hat{F}\|_{op}C(n,m)$, where $C(n,m):=\sum_{M=n}^{m-1}[c^{+}_{M}]^{-1}$ and $q_{0}$ is a constant. In particular, for $0\leq b<a\leq 1$, if we take $J\sim N^{a}$ and $|m-n|\sim N^{b}$, then we have $|\langle J,n|\hat{F}|J,n\rangle-\langle J,m|\hat{F}|J,m\rangle|=\|\hat{F}\|_{op}\cdot O(dN^{-(a-b)})$.
@@
We can obtain a bound on $C(n,m)$ as follows. Consider the function $g(x):=[(J-x)(J+x+1)]^{-1/2}$ in the domain $x\in(-J-1,J)$. It is easy to see that the minimum of $g(x)$ is at $x^{*}=-1/2$, and $g(x)$ is a decreasing function from x=J1x=-J-1 to $x=x^{*}$ and an increasing function from $x=x^{*}$ to $x=J$. It is also easy to check that $g(x)=g(2x^{*}-x)$. We first consider the case $n<0<m$. We separate $C(n,m)$ into two pieces, $C(n,m)=C_{1}(n)+C_{2}(m)$, where $C_{1}(n):=\sum_{\ell=n}^{\ell=-1}[c^{+}_{\ell}]^{-1}$ and $C_{2}(m):=\sum_{\ell=0}^{m-1}[c^{+}_{\ell}]^{-1}$. These two pieces can be bounded as
$\int_{n}^{0}g(x)dx$ $<C_{1}(n)<\int_{n}^{0}g(x-1)dx~,$
$\int_{0}^{m}g(x-1)dx$ $<C_{2}(m)<\int_{0}^{m}g(x)dx~.$ (S8)
The integrals of the upper bounds are given by $\int_{0}^{m}g(x)dx=2\arcsin\left(\sqrt{\frac{J}{2J+1}}\right)-2\arcsin\left(\sqrt{\frac{J-m}{2J+1}}\right)$ and $\int_{n}^{0}dxg(x-1)=\int_{n}^{0}dxg(-x)=\int_{0}^{-n}dyg(y)=2\arcsin\left(\sqrt{\frac{J}{2J+1}}\right)-2\arcsin\left(\sqrt{\frac{J+n}{2J+1}}\right)$.
@@
If $0<n<m$, then we only need to bound $\sum_{\ell=n}^{m-1}[c^{+}_{\ell}]^{-1}<\int_{n}^{m}g(x)dx=\int_{0}^{m}g(x)dx-\int_{0}^{n}g(x)dx=|m-n|J^{-1}+O(|m-n|J^{-2})$; if $n<m<0$, we bound $\sum_{\ell=n}^{m-1}[c^{+}_{\ell}]^{-1}<\int_{n}^{m}g(x-1)dx=\int_{-m}^{-n}g(x)dx=|m-n|J^{-1}+O(|m-n|J^{-2})$.
Substituting the above results to Eq. (S3), we have
$|\langle J,{m}|\hat{F}|J,{m}\rangle-\langle J,n|\hat{F}|J,n\rangle|\leq\|\hat{F}\|_{op}\cdot\left(\frac{dq_{0}|m-n|}{J}+O\left(\frac{d|m-n|}{J^{2}}\right)\right)~.$ (S12)
@@
$|\langle J,{m}|\hat{F}|J,{m}\rangle-\langle J,n|\hat{F}|J,n\rangle|$ $=\|\hat{F}\|_{op}\cdot[O(dN^{b-a})+O(dN^{b-2a})]$
$=\|\hat{F}\|_{op}\cdot O(dN^{b-a})~.$ (S13)
@@
Theorem 1: Consider the code space $\mathfrak{C}=\text{span}\{|J\!,\!M\rangle;M\!=\!M_{\mathrm{min}},M_{\mathrm{min}}\!+\!\Delta,M_{\mathrm{min}}\!+\!2\Delta,\dots,M_{\mathrm{max}}\}$, where $J\!\sim\!N^{a}$, $M_{\mathrm{max}}\!-\!M_{\mathrm{min}}\!\sim\!N^{b}$, and $b\!<\!a\!\leq 1$. The code $\mathfrak{C}$ forms an $(\!(N,O[b\log_{2}N])\!)$ AQECC against $d$-local noise on known sites, with inaccuracy $\epsilon(d)=O[\sqrt{d}(2s+1)^{d}N^{-(a-b)/2}]$ if $\Delta\geq 2sd+1$.